OpenStudy (anonymous):

Given that p and 2 are the roots of the quadratic equation q(1-x)=(x+3)(1-2x),where q is a constant,calculate the values of p and of q. @rational @radar

OpenStudy (anonymous):

\[q-qx=2x^2-x+3-6x\]\[-2x^2+7x-qx+q-3=0\]

OpenStudy (anonymous):

\[(x+3)(1-2x)=x-2x^2+3-6x\]

OpenStudy (anonymous):

ops,thnx @mustafa2014

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

\[2x^2+5x-qx+q-3=0\]

OpenStudy (anonymous):

divide both side by 2

OpenStudy (anonymous):

q-qx=x-2x^2+3-6x q-qx-x+2x^2-3+6x=0 x^2+(5-q)/2x+(q-3)/2=0 -2-p=(5-q)/2 -4-2p=5-q q-2p=9 2p=(q-3)/2 4p-q=-3 4p-9-2p=-3 2p=6 p=3 q=15 x^2-5x+6=0

OpenStudy (anonymous):

i don't get it. why -2-p?

OpenStudy (anonymous):

cause x^2+(5-q)/2x+(q-3)/2=0 -2-p=(5-q)/2 -4-2p=5-q q-2p=9 2p=(q-3)/2

OpenStudy (anonymous):

p and 2 are the roots means (x-p)(x-2)=0 x^2-2x-px+2p=0

OpenStudy (anonymous):

x^2-(2+p)x+2p=0 and we have x^2+(5-q)/2x+(q-3)/2=0

OpenStudy (anonymous):

Thnx for ur explanation @mustafa2014 :)

OpenStudy (anonymous):

np :)